# Project Page: https://lizhiqi49.github.io/MVControl/

### v2

@@ Abstract
While text-to-3D and image-to-3D generation tasks have received considerable attention, one important but under-explored field between them is controllable text-to-3D generation, which we mainly focus on in this work. To address this task, 1) we introduce Multi-view ControlNet (MVControl), a novel neural network architecture designed to enhance existing pre-trained multi-view diffusion models by integrating additional input conditions, such as edge, depth, normal, and scribble maps. Our innovation lies in the introduction of a conditioning module that controls the base diffusion model using both local and global embeddings, which are computed from the input condition images and camera poses. Once trained, MVControl is able to offer 3D diffusion guidance for optimization-based 3D generation. And, 2) we propose an efficient multi-stage 3D generation pipeline that leverages the benefits of recent large reconstruction models and score distillation algorithm. Building upon our MVControl architecture, we employ a unique hybrid diffusion guidance method to direct the optimization process. In pursuit of efficiency, we adopt 3D Gaussians as our representation instead of the commonly used implicit representations. We also pioneer the use of SuGaR, a hybrid representation that binds Gaussians to mesh triangle faces. This approach alleviates the issue of poor geometry in 3D Gaussians and enables the direct sculpting of fine-grained geometry on the mesh. Extensive experiments demonstrate that our method achieves robust generalization and enables the controllable generation of high-quality 3D content.
Keywords:
Controllable 3D Generation Gaussian Splatting SuGaR
00footnotetext: $\dagger$ Corresponding author.
@@ 2 Related Work
Multi-view Diffusion Models. The success of text-to-image generation via large diffusion models inspires the development of multi-view image generation. Commonly adopted approach is to condition on a diffusion model by an additional input image and target pose [31, 34, 32] . Different from those methods, Chan et al. recently proposes to learn 3D scene representation from a single or multiple input images and then exploit a diffusion model for target novel view image synthesis [10]. Instead of generating a single target view image, MVDiffusion [58] proposes to generate multi-view consistent images in one feed-forward pass. They build upon a pre-trained diffusion model to have better generalization capability. MVDream [50] introduces a method for generating consistent multi-view images from a text prompt. They achieve this by fine-tuning a pre-trained diffusion model using a 3D dataset. The trained model is then utilized as a 3D prior to optimize the 3D representation through Score Distillation Sampling. Similar work ImageDream[60] substitutes the text condition with an image. While prior works can generate impressive novel/multi-view consistent images, fine-grained control over the generated text-to-multi-view images is still difficult to achieve, as what ControlNet [69] has achieved for text-to-image generation. Therefore, we propose a multi-view ControlNet (i.e. MVControl) in this work to further advance diffusion-based multi-view image generation.
@@
3D Generation Tasks. The exploration of generating 3D models can typically be categorized into two approaches. The first is the SDS-based optimization method, initially proposed by DreamFusion[42], which aims to extract knowledge for 3D generation through the utilization of pre-trained large image models. SDS-based method benefits from not requiring expansive 3D datasets and has therefore been extensively explored in subsequent works[29, 11, 63, 52, 56, 70, 66, 43]. These works provide insights of developing more sophisticated score distillation loss functions [63, 43, 52], refining optimization strategies [70, 29, 11, 52, 56], and employing better 3D representations [11, 63, 52, 56, 66], thereby further enhancing the quality of the generation. Despite the success achieved by these methods in generating high-fidelity 3D assets, they usually require hours to complete the text-to-3D generation process. In the contrary, feed-forward 3D native methods can produce 3D assets within seconds after training on extensive 3D datasets[14]. Researchers have explored various 3D representations to achieve improved results, such as volumetric representation[6, 15, 64, 28], triangular mesh[54, 16, 13, 67], point cloud[2, 1], implicit neural representation[41, 36, 12, 48, 9, 62, 26, 19], as well as the recent 3D Gaussian[55]. Some of these methods can efficiently generate 3D models that fully satisfy the input conditions. While some methods efficiently generate 3D models that meet input conditions, 3D generative methods, unlike image generative modeling, struggle due to limited 3D training assets. This scarcity hinders their ability to produce high-fidelity and diverse 3D objects. Our method merges both approaches: generating a coarse 3D object with a feed-forward method conditioned on MVControl’s output, then refining it using SDS loss for the final representation.
@@
Optimization based Mesh Generation. The current single-stage mesh generation method, such as MeshDiffusion[33], struggles to produce high-quality mesh due to its highly complex structures. To achieve high grade mesh in both geometry and texture, researchers often turn to multi-stage optimization-based methods[29, 11, 52]. These methods commonly use non-mesh intermediate representations that are easy to process, before transforming them back into meshes with mesh reconstruction methods, which can consume a long optimization time. DreamGaussian[56] refer to a more efficient representation, 3D Gaussian, to effectively reduce the training time. However, extracting meshes from millions of unorganized tiny 3D Gaussians remains challenging. LGM[55] presents a new mesh extraction method for 3D Gaussian but still relies on implicit representation. In contrast, we adopts a fully explicit representation, a hybrid of mesh and 3D Gaussian as proposed by SuGaR[17]. This approach enables us to achieve high-quality mesh generation within reasonable optimization time.
@@ 3.1 Preliminary
ControlNet. ControlNet [69] facilitates pretrained large text-to-image diffusion models to accommodate additional input conditions (e.g., canny edges, sketches, depth maps, etc.) alongside text prompts, thereby enabling precise control over generated content. It directly copies the structure and weights of SD’s encoder blocks and mid block as an additional module into the backbone model. The output feature maps of each layer of the ControlNet module’s encoder blocks and mid block are injected into the corresponding symmetric layers of the backbone model’s decoder blocks and mid block using 1x1 convolutions. To preserve the generation capability of the backbone model and ensure smooth training initiation, the 1x1 convolutions are initialized to zero.
@@
Figure 2: Overview of proposed method. (a) MVControl consists of a frozen multi-view diffusion model and a trainable MVControl. (b) Our model takes care of all input conditions to control the generation process both locally and globally through a conditioning module. (c) Once MVControl is trained, we can exploit it to serve a hybrid diffusion prior for controllable text-to-3D content generation via SDS optimization procedure.
@@ 3.2 Multi-view ControlNet
Inspired by ControlNet in controlled text-to-image generation and recently released text-to-multi-view image diffusion model (e.g. MVDream), we aim to design a multi-view version of ControlNet (i.e. MVControl) to achieve controlled text-to-multi-view generation. As shown in Fig. 2, we follow similar architecture style as ControlNet, i.e. a locked pre-trained MVDream and a trainable control network. The main insight is to preserve the learned prior knowledge of MVDream, while train the control network to learn the inductive bias with small amount of data. The control network consists of a conditioning module and a copy of the encoder network of MVDream. Our main contribution lies at the conditioning module and we will detail it as follows.
@@
While MVDream is pretrained with absolute camera poses, the conditioning module exploit relative poses as input. We experimentally find that the network hardly converges due to the mismatch of both coordinate frames. We therefore exploit an additional network $\mathcal{M}_{2}$ to learn the transformation and output a global embedding $e^{g}_{t,c,v_{*}}$ to replace the original camera matrix embedding of MVDream and add on timestep embeddings of both MVDream and MVControl part, so that semantical and view-dependent features are injected globally.
@@ 3.3 Controllable 3D Textured Mesh Generation
$\nabla_{\theta}\mathcal{L}_{SDS}^{hybrid}=\lambda_{2D}\nabla_{\theta}\mathcal{L}_{SDS}^{2D}(\mathbf{x}_{r}=g(\theta,\mathcal{V}_{r});t,y)+\lambda_{3D}\nabla_{\theta}\mathcal{L}_{SDS}^{3D}(\mathbf{x}_{*}=g(\theta,\mathcal{V}_{*});t,y,h),$ (2)
where $\lambda_{1}$ and $\lambda_{2}$ are the strength of 2D and 3D prior respectively. To enhance the learning of geometry during the Gaussians optimization stage, we employ a Gaussian rasterization engine capable of rendering depth and alpha values [5]. Specifically, in addition to color images, depth $\hat{d}$ and alpha $\hat{m}$ of the scene are also rendered, and we estimate the surface normal $\hat{n}$ by taking the derivative of $\hat{d}$. Consequently, the total variation (TV) regularization terms [46] on these components, denoted as $\mathcal{L}_{TV}^{d}$ and $\mathcal{L}_{TV}^{n}$, are calculated and incorporated into the hybrid SDS loss. Furthermore, as the input conditions are invariably derived from existing images, a foreground mask $m_{gt}$ is generated during the intermediate process. Therefore, we compute the mask loss $\mathcal{L}_{mask}=\text{MSE}(\hat{m},m{gt})$ to ensure the sparsity of the scene. Thus, the total loss for Gaussian optimization is expressed as:
@@ 4 Experiments
4.1 Implementation Details
Training Data. We employ the multi-view renderings from the publicly available large 3D dataset, Objaverse [14], to train our MVControl. Initially, we preprocess the dataset by removing all samples with a CLIP-score lower than 22, based on the labeling criteria from [53]. This filtering results in approximately 400k remaining samples.Instead of utilizing the names and tags of the 3D assets, we employ the captions from [35] as text descriptions for our retained objects. Following the approach in [50], our network is trained on both 2D and 3D datasets. Specifically, we randomly sample images from the AES v2 subset of LAION [47] with a 30% probability during training to ensure the network retains its learned 2D image priors. Additionally, we sample from our curated 3D/multi-view image dataset with a 70% probability to learn 3D knowledge. The process of preparing different types of condition images will be detailed in our appendix.
Training details of MVControl. We initialize our network by leveraging the weights of the pretrained MVDream and ControlNet. All connections between the locked and trainable networks are initialized to zero. MVControl is trained at a resolution of 256, with a chosen batch size of 2560 images. The model undergoes fine-tuning for 50,000 steps using a conservative learning rate of $4e^{-5}$, employing the AdamW optimizer [25] on 8 Nvidia Tesla A100 GPUs. Similar to the approach in [69], we randomly drop the text prompt as empty with a 50% chance during training to facilitate classifier-free learning and enhance the model’s understanding of input condition images.
3D Generation. The coarse Gaussian optimization stage comprises a total of 3000 steps. During the initial 1500 steps, we perform straightforward 3D Gaussian optimization with split and prune. After step 1500, we cease Gaussian split and prune and instead introduce SuGaR [17] regularization terms to refine the scene. We set the number of nearest neighbors to 16, updating them every 50 steps. During training’s final phase, we prune all Gaussians with opacity below $\bar{\sigma}=0.5$. We use $\lambda_{2D}=0.1$ and $\lambda_{3D}=0.01$ for 2D and 3D diffusion guidance, with resolutions of 512 and 256 respectively. For coarse SuGaR extraction, we target 2e5 vertices, binding 6 Gaussians to each triangle. In the SuGaR refinement stage, we optimize the scene for 5000 steps with VSD loss [63].
@@ 4.2 Qualitative Comparisons
Multi-view Image Generation. To assess the controllability of our MVControl, we conduct experiments on MVDream both with and without MVControl attached. In the first case, MVDream fails to generate the correct contents according to the given prompt, producing a standing cat without clothes, which contradicts the prompt. In contrast, with the assistance of MVControl, it successfully generates the correct contents. The second case also demonstrates that our MVControl effectively controls the generation of MVDream, resulting in highly view-consistent multi-view images.
Figure 5: Comparison with baseline 3D generation methods. Our method yields more delicate texture, and generates much better meshes than the compared methods.
Figure 6: Comparison with implicit representation based 3D generation methods. We compare our model with the state-of-the-art image-to-3D method DreamCraft3D [52] and our base model MVDream [50].
3D Gaussian-based Mesh Generation. Given that our 3D generation pipeline aims to produce textured mesh from 3D Gaussians, we compare our method with recent Gaussian-based mesh generation approaches, DreamGaussian [56] and LGM [55], both of which can be conditioned on RGB images. For a fair comparison, we generate a 2D RGB image using the 2D ControlNet as the condition for the compared methods. As illustrated in Fig. 5, DreamGaussian struggles to generate the geometry for most of the examples, resulting in many broken and hollow areas in the generated meshes. LGM performs better than DreamGaussian, however, its extracted meshes lack details and still contain broken areas in some cases. In contrast, our method produces fine-grain meshes with more delicate textures, even without an RGB condition. Due to space limitations, the textual prompts are not provided in Fig. 5, and we will include them in our appendix. All the images in the figure are rendered using Blender.
Implicit Representation-based Mesh Generation. To provide a comprehensive evaluation of our model, we extend the application of MVControl to 3D generation tasks based on implicit representations. We compare our approach with the state-of-the-art image-to-3D method, DreamCraft3D [52]. We achieve this by employing a straightforward coarse-to-fine optimization procedure. Initially, we generate a coarse NeuS [59] model, followed by its transfer to a coarse DMTet [49] for further geometry optimization. The final stage focuses on texture refinement, with the geometry fixed. The hybrid diffusion guidance is utilized in all three stages, with varying weight terms. As illustrated in Fig. 6, our method is capable of generating 3D assets with geometry and texture comparable to DreamCraft3D, even in the absence of RGB signals. Furthermore, we conducted experiments on our base model, MVDream, using the same prompt. The results demonstrate that the prompt alone cannot precisely control the generation of MVDream (evidenced by the sitting buddha conflicting with the prompt) without the use of our MVControl.
@@ 4.3 Quantitative Comparisons
In this section, we adopt CLIP-score [39] to evaluate both the compared methods and our method. Given that our task involves two conditions, namely the prompt and the condition image of the reference view, we calculate both image-text and image-image similarities. For each object, we uniformly render 36 surrounding views. The image-text similarity, denoted as CLIP-T, is computed by averaging
@@
DreamGaussian [56] 0.200 0.847
LGM [55] 0.228 0.872
MVControl(Ours) 0.245 0.897
Table 1: Quantitative Comparison. We calculate both image-text and image-image similarity with CLIP score.
the similarities between each view and the given prompt.Similarly, the image-image similarity, referred to as CLIP-I, is the mean similarity between each view and the reference view. The results, calculated for a set of 60 objects, are reported in Table 1. When employing our method, the condition type used for each object is randomly sampled from edge, depth, normal, and scribble map. Additionally, the RGB images for DreamGaussian and LGM are generated using 2D ControlNet with the same condition image and prompt.
Figure 7: Ablation study on conditioning module.
Figure 7: Ablation study on conditioning module.
Figure 8: Ablation study on Gaussian optimization stage.
4.4 Ablation Study
@@
Hybrid Diffusion Guidance. We conduct ablation study on hybrid diffusion guidance utilized in the Gaussian optimization stage. As illustrated in Fig. 8 (top right), when excluding $\nabla_{\theta}\mathcal{L}_{SDS}^{3D}$ provided by our MVControl, the generated 3D Gaussians lack texture details described in the given condition edge map. For instance, the face of the rabbit appears significantly blurrier without $\nabla_{\theta}\mathcal{L}_{SDS}^{3D}$.
Losses on Rendered Normal Maps. The normal-related losses in our method are alternately calculated using SDS loss with the normal map as input in stage 2, and the normal TV regularization term. We conduct experiments by dropping all of them in stage 2, and the results are illustrated in Fig. 8 (bottom left). Compared to our full method, the surface normal of 3D Gaussians deteriorates without the normal-related losses.
Multi-stage Optimization. We also assess the impact of different optimization stages, as shown in Fig. 9. Initially, in stage 1, the coarse Gaussians exhibit poor geometry consistency. However, after the Gaussian optimization stage, they become view-consistent, albeit with blurry texture. Finally, in the SuGaR refinement stage, the texture of the 3D model becomes fine-grained and of high quality.

### v3

# Controllable Text-to-3D Generation via Surface-Aligned Gaussian Splatting

@@ Abstract
While text-to-3D and image-to-3D generation tasks have received considerable attention, one important but under-explored field between them is controllable text-to-3D generation, which we mainly focus on in this work. To address this task, 1) we introduce Multi-view ControlNet (MVControl), a novel neural network architecture designed to enhance existing pre-trained multi-view diffusion models by integrating additional input conditions, such as edge, depth, normal, and scribble maps. Our innovation lies in the introduction of a conditioning module that controls the base diffusion model using both local and global embeddings, which are computed from the input condition images and camera poses. Once trained, MVControl is able to offer 3D diffusion guidance for optimization-based 3D generation. And, 2) we propose an efficient multi-stage 3D generation pipeline that leverages the benefits of recent large reconstruction models and score distillation algorithm. Building upon our MVControl architecture, we employ a unique hybrid diffusion guidance method to direct the optimization process. In pursuit of efficiency, we adopt 3D Gaussians as our representation instead of the commonly used implicit representations. We also pioneer the use of SuGaR, a hybrid representation that binds Gaussians to mesh triangle faces. This approach alleviates the issue of poor geometry in 3D Gaussians and enables the direct sculpting of fine-grained geometry on the mesh. Extensive experiments demonstrate that our method achieves robust generalization and enables the controllable generation of high-quality 3D content. The source code is available at our website: https://lizhiqi49.github.io/MVControl.
footnotetext: Corresponding author.
@@ 1 Introduction
Figure 2: Architecture of proposed MVControl. (a) MVControl consists of a frozen multi-view diffusion model and a trainable MVControl. (b) Our model takes care of all input conditions to control the generation process both locally and globally through a conditioning module. (c) Once MVControl is trained, we can exploit it to serve a hybrid diffusion prior for controllable text-to-3D content generation via SDS optimization procedure.
@@ 2 Related Work
Multi-view Diffusion Models. The success of text-to-image generation via large diffusion models inspires the development of multi-view image generation. Commonly adopted approach is to condition on a diffusion model by an additional input image and target pose [30, 33, 31]. Unlike those methods, Chan et al. recently proposed to learn 3D scene representation from a single or multiple input images and then exploit a diffusion model for target novel view image synthesis [10]. Instead of generating a single target view image, MVDiffusion [57] proposes to generate multi-view consistent images in one feed-forward pass. They build upon a pre-trained diffusion model to have better generalization capability. MVDream [49] introduces a method for generating consistent multi-view images from a text prompt. They achieve this by fine-tuning a pre-trained diffusion model using a 3D dataset. The trained model is then utilized as a 3D prior to optimize the 3D representation through Score Distillation Sampling. Similar work ImageDream[58] substitutes the text condition with an image. While prior works can generate impressive novel/multi-view consistent images, fine-grained control over the generated text-to-multi-view images is still difficult to achieve, as what ControlNet [68] has achieved for text-to-image generation. Therefore, we propose a multi-view ControlNet (i.e. MVControl) in this work to further advance diffusion-based multi-view image generation.
@@
3D Generation Tasks. The exploration of generating 3D models can typically be categorized into two approaches. The first is SDS-based optimization method, initially proposed by DreamFusion[41], which aims to extract knowledge for 3D generation through the utilization of pre-trained large image models. SDS-based method benefits from not requiring expansive 3D datasets and has therefore been extensively explored in subsequent works[28, 11, 62, 51, 54, 69, 65, 42]. These works provide insights of developing more sophisticated score distillation loss functions [62, 42, 51], refining optimization strategies [69, 28, 11, 51, 54], and employing better 3D representations [11, 62, 51, 54, 65], thereby further enhancing the quality of the generation. Despite the success achieved by these methods in generating high-fidelity 3D assets, they usually require hours to complete the text-to-3D generation process. On the contrary, feed-forward 3D native methods can produce 3D assets within seconds after training on extensive 3D datasets[14]. Researchers have explored various 3D representations to achieve improved results, such as volumetric representation[6, 15, 63, 27], triangular mesh[53, 16, 13, 66], point cloud[2, 1], implicit neural representation[40, 35, 12, 47, 9, 61, 25, 19], as well as the recent 3D Gaussian[56]. While some methods efficiently generate 3D models that meet input conditions, 3D generative methods, unlike image generative modeling, struggle due to limited 3D training assets. This scarcity hinders their ability to produce high-fidelity and diverse 3D objects. Our method merges both approaches: generating a coarse 3D object with a feed-forward method conditioned on MVControl’s output, then refining it using SDS loss for the final representation.
@@
Optimization-based Mesh Generation. The current single-stage mesh generation method, such as MeshDiffusion[32], struggles to produce high-quality mesh due to its highly complex structures. To achieve high grade mesh in both geometry and texture, researchers often turn to multi-stage optimization-based methods[28, 11, 51]. These methods commonly use non-mesh intermediate representations that are easy to process, before transforming them back into meshes with mesh reconstruction methods, which can consume a long optimization time. DreamGaussian[54] refer to a more efficient representation, 3D Gaussians, to effectively reduce the training time. However, extracting meshes from millions of unorganized tiny 3D Gaussians remains challenging. LGM[56] presents a new mesh extraction method for 3D Gaussians but still relies on implicit representation. In contrast, we adopt a fully explicit representation, a hybrid of mesh and 3D Gaussians as proposed by SuGaR[17]. This approach enables us to achieve high-quality mesh generation within reasonable optimization time.
@@ 3.2 Multi-view ControlNet
Inspired by ControlNet in controlled text-to-image generation and recently released text-to-multi-view image diffusion model (e.g. MVDream), we aim to design a multi-view version of ControlNet (i.e. MVControl) to achieve controlled text-to-multi-view generation. As shown in Fig. 2, we follow similar architecture style as ControlNet, i.e. a locked pre-trained MVDream and a trainable control network. The main insight is to preserve the learned prior knowledge of MVDream, while training the control network to learn the inductive bias with a small amount of data. The control network consists of a conditioning module and a copy of the encoder network of MVDream. Our main contribution lies at the conditioning module and we will detail it as follows.
@@
While MVDream is pretrained with absolute camera poses, the conditioning module exploits relative poses as input. We experimentally find that the network hardly converges due to the mismatch of both coordinate frames. We therefore exploit an additional network $\mathcal{M}_{2}$ to learn the transformation and output a global embedding $e^{g}_{t,c,v_{*}}$ to replace the original camera matrix embedding of MVDream and add on timestep embeddings of both MVDream and MVControl part, so that semantical and view-dependent features are injected globally.
Figure 4: Comparison with baseline 3D generation methods. Our method yields more delicate texture, and generates much better meshes than the compared methods. We use different color blocks to emphasize that our method only takes the conditioning image rather than RGB as input. Corresponding textual prompts are provided in appendix.
@@ 3.3 Controllable 3D Textured Mesh Generation
$\nabla_{\theta}\mathcal{L}_{SDS}^{hybrid}=\lambda_{2D}\nabla_{\theta}\mathcal{L}_{SDS}^{2D}(\mathbf{x}_{r}=g(\theta,\mathcal{V}_{r});t,y)$ (2)
$+\lambda_{3D}\nabla_{\theta}\mathcal{L}_{SDS}^{3D}(\mathbf{x}_{*}=g(\theta,\mathcal{V}_{*});t,y,h),$
where $\lambda_{1}$ and $\lambda_{2}$ are the strength of 2D and 3D prior respectively. To enhance the learning of geometry during the Gaussians optimization stage, we employ a Gaussian rasterization engine capable of rendering depth and alpha values [5]. Specifically, in addition to color images, depth $\hat{d}$ and alpha $\hat{m}$ of the scene are also rendered, and we estimate the surface normal $\hat{n}$ by taking the derivative of $\hat{d}$. Consequently, the total variation (TV) regularization terms [45] on these components, denoted as $\mathcal{L}_{TV}^{d}$ and $\mathcal{L}_{TV}^{n}$, are calculated and incorporated into the hybrid SDS loss. Furthermore, as the input conditions are invariably derived from existing images, a foreground mask $m_{gt}$ is generated during the intermediate process. Therefore, we compute the mask loss $\mathcal{L}_{mask}=\text{MSE}(\hat{m},m_{gt})$ to ensure the sparsity of the scene. Thus, the total loss for Gaussian optimization is expressed as:
@@ 4.1 Qualitative Comparisons
Multi-view Image Generation. To assess the controlling capacity of our MVControl, we conduct experiments on MVDream both with and without MVControl attached as shown in Fig. 5. In the first case, MVDream fails to generate the correct contents according to the given prompt, producing a squatting cat without clothes, which contradicts the prompt. In contrast, it successfully generates the correct contents with the assistance of MVControl. The second case also demonstrates that our MVControl effectively controls the generation of MVDream, resulting in highly view-consistent multi-view images.
3D Gaussian-based Mesh Generation. Given that our 3D generation pipeline aims to produce textured mesh from 3D Gaussians, we compare our method with recent Gaussian-based mesh generation approaches, DreamGaussian [54] and LGM [56], both of which can be conditioned on RGB images. Moreover, we also take the state-of-the-art image-to-3D generation method, DreamCraft3D [51] into comparison. For fair comparison, we generate 2D RGB images without cherry picking using the pre-trained 2D ControlNet as the input for the compared methods. As illustrated in Fig. 4, DreamGaussian struggles to generate the geometry for most of the examples, resulting in many broken and hollow areas in the generated meshes. LGM performs better than DreamGaussian, however, its extracted meshes lack details and still contain broken areas in some cases. Although DreamCraft3D can produce unbroken shapes, it still suffer from unsmoothed surface in the meshes. In contrast, our method produces fine-grain meshes with more delicate textures, even without an RGB condition. Due to space limitations, the textual prompts are not provided in Fig. 4, and we will include them in our appendix.
[…]
Table 1: Quantitative comparison with baselines. Our method achieves the best result.
@@ 4.2 Quantitative Comparisons
In this section, we adopt CLIP-score [38] to evaluate the compared methods and our method. We calculate both image-text and image-image similarities. For each object, we uniformly render 36 surrounding views. The image-text similarity, denoted as CLIP-T, is computed by averaging the similarities between each view and the given prompt.Similarly, the image-image similarity, referred to as CLIP-I, is the mean similarity between each view and the reference view. The results, calculated for a set of 60 objects, are reported in Table 1. When employing our method, the condition type used for each object is randomly sampled from edge, depth, normal, and scribble map. Additionally, the RGB images for DreamGaussian and LGM are generated using 2D ControlNet with the same condition image and prompt. Our method achieves the best performance in terms of both the CLIP-T and CLIP-I score.
4.3 Ablation Study
@@
Stage 2 w/o $\nabla_{\theta}\mathcal{L}_{SDS}^{3D}$ 0.245 0.866
Stage 2 w/o normal losses 0.263 0.876
Full stage 2 0.267 0.882
Stage 1 only 0.230 0.859
Full method 0.279 0.909
Table 2: Quantitative comparison of ablation study.
Figure 6: Ablation study on conditioning module. It achieves precise controlling on multi-view image generation with our conditioning module.
@@
Hybrid Diffusion Guidance. We conduct ablation study on hybrid diffusion guidance utilized in the Gaussian optimization stage. As illustrated in Fig. 7 (top right), when excluding $\nabla_{\theta}\mathcal{L}_{SDS}^{3D}$ provided by our MVControl, the generated 3D Gaussians lack texture details described in the given condition edge map. For instance, the face of the rabbit appears significantly blurrier without $\nabla_{\theta}\mathcal{L}_{SDS}^{3D}$. The quantitative evaluation is provided in Table 2 (line 1 and 3).
Losses on Rendered Normal Maps. The normal-related losses in our method are alternately calculated using SDS loss with the normal map as input in stage 2, and the normal TV regularization term. We conduct experiments by dropping all of them in stage 2, and the results are illustrated in Fig. 7 (bottom left). Compared to our full method, the surface normal of 3D Gaussians deteriorates without the normal-related losses. The corresponding quantitative results are provided in Table 2 (line 2 and 3).
Figure 7: Ablation study on Gaussian optimization stage. The qualitative results demonstrate the effects of the $\nabla_{\theta}\mathcal{L}_{SDS}^{3D}$ and normal related loss terms during optimization.
Multi-stage Optimization. We also assess the impact of different optimization stages, as shown in Fig. 8. Initially, in stage 1, the coarse Gaussians exhibit poor geometry consistency. However, after the Gaussian optimization stage, they become view-consistent, albeit with blurry texture. Finally, in the SuGaR refinement stage, the texture of the 3D model becomes fine-grained and of high quality. We have also provided the quantitative evaluation on different optimization stages in Table 2 (the lower 3 lines).